# Unlocking the power of global quantum gates with machine learning

### v1

@@ Abstract
In conventional circuit-based quantum computing architectures, the standard gate set includes arbitrary single-qubit rotations and two-qubit entangling gates. However, this choice is not always aligned with the native operations available in certain hardware, where the natural entangling gates are not restricted to two qubits but can act on multiple, or even all, qubits simultaneously. However, leveraging the capabilities of global quantum operations for algorithm implementations is highly challenging, as directly compiling local gate sequences into global gates usually gives rise to a quantum circuit that is more complex than the original one. Here, we circumvent this difficulty using a variational approach. Specifically, we propose a parameterized circuit ansatze composed of a finite number of global gates and layers of single-qubit unitaries, which can be implemented in constant time. Furthermore, by construction, these ansatze are equivalent to linear depth local-gate quantum circuits and are highly expressible. We demonstrate the capabilities of this approach by applying it to the problem of ground state preparation for the Heisenberg model and the toric code Hamiltonian, highlighting its potential to offer a practical quantum advantage.
@@
In this work, we address this gap with a variational approach (Figure 1). Specifically, we implement variational quantum algorithms using circuit ansatzes that are directly constructed from alternating layers of global gates and single-qubit rotations. This design restricts the circuit to a constant number of global gates, ensuring that the implementations remain low-cost and are suitable for near-term quantum computers supporting such gates. Despite these constraints, we demonstrate that the ansatzes are capable of preparing a diverse range of quantum states, including those with long-range entanglement.
@@ I Global Gates
Global gates are unitary operations that simultaneously act on multiple or all qubits in a system. They are native to several platforms and have been demonstrated experimentally across a range of hardware [9, 10, 11, 5, 12, 13, 14]. They are typically achieved by driving the systems through their native underlying interactions and with global control pulses. For example, in Rydberg atom systems, global operations can be implemented using resonant laser pulses over a static atomic arrangement [14]. Superconducting systems can emulate global gates by coupling the Josephson junction qubits to a shared resonator mode [16].
@@
In addition, as single-qubit gates can be implemented efficiently in high fidelity with small execution time compared to the entangling operations on the physical device, we treat single-qubit gates a free resource. This also makes working with a variety of entangling gates equivalent so long as they can be obtained from the GCZ gates using a simple single-qubit basis change. For instance, we treat the GCX gate
@@ II.1 Barren Plateau
One key requirement for any successful ansatz is its ability to train well, starting for any arbitrary initial state and for a wide variety of Hamiltonians. The gradient-based optimizers rely on sufficiently large gradient values so that the parameters of the ansatz can be updated. However, it is well known that sufficiently complex ansatze (as random as a unitary $2$-design) admit the infamous “barren plateau” phenomenon [20, 21]. That is, the variance of the gradient of the cost function, ${\rm Var}\left[\partial_{\vec{\theta}}C\right]$ decays exponentially with the system size, making the landscape of the cost function essentially flat and untrainable.
Numerous methods have been adopted to surpass the barren plateau issue, with various benefits and drawbacks. Here, we focus on a recently proposed solution named the Finite Local-Depth Circuits (FLDC) [22]. These circuits are specifically structured such that each qubit encounters only a finite number of quantum gates (hence local depth), while the total depth of the circuit remains linear in the system size. FLDC were demonstrated to be free of barren plateau, sufficiently expressive to represent even long-range entangled states, and hard to classically simulate efficiently. These features strike a balance between maintaining rich entanglement and preserving a non-vanishing gradient landscape. More importantly, FLDC can be naturally adopted to incorporate global quantum gates, as will be discussed in the next section.
@@ II.2 Global Gate Ansatze
Our general strategy for constructing the global gate ansatze starts from building the circuit with a finite number ($k$) of alternating layers of single-qubit gates and two-qubit gates. See Figure 2 for an illustration. The overall depth of the circuit—with respect to local gates—is linear in the number of qubits. However, the two-qubit layers are designed to have finite local depth and are therefore guaranteed to be free of barren plateau. Crucially, each layer of the two-qubit gates consists of only the same type of gates, i.e., ${\rm CZ}(\theta)$ as defined in (2), such that they can be implemented using a single global gate GCZ in equation (1).
We compare a variety of ansatze as constructed with the above general procedure. The detailed structures of the ansatze also depend on the underlying architecture (qubit connectivity). For simplicity, we first describe ansatze with qubits arranged in a $1$-D geometry. In this case, each layer of 2-qubit unitaries connects only neighboring sites in a ladder format (Figure 2).
@@
  • Cartan Ansatz: We also compare the above ansatzes, which can be implemented with a finite number of global gates, to the most general FLDC, where each two-qubit gate can be an arbitrary two-qubit unitary. The name is inspired from the Cartan decomposition of arbitrary 2-qubit unitaries. This ansatz cannot be efficiently realized using global gates but nevertheless serves as a benchmark for our global gate circuits.
Note that the ansatze constructed using the above procedure has finite local depth in terms of local gate count and is, therefore, barren plateau free. We illustrate this fact with numerical simulations in Figure 3 (a,b).
@@ IV Discussion
In this work, we developed a variational approach to harness the power of global quantum gates, native to many quantum computing platforms. Our proposed circuit ansatze consists of a finite number of global quantum gates interspersed with single-qubit rotations, enabling efficient implementation within finite time. These circuits are as powerful as linear-depth circuits composed of local quantum gates and, in the meanwhile, avoid the barren plateau problem, making them efficiently trainable. We demonstrated the efficacy of this approach by using the global gate ansatze to prepare the ground states of the Heisenberg model and the Toric code model, achieving fast convergence and high accuracy. Future work could explore extending this method to other applications, such as unitary compiling.
Although we focused on the global control-Z gate, our approach can naturally generalize to other sets of global operations. Another promising direction is to investigate scenarios where the two-qubit layers have logarithmic local depth rather than the finite local depth arrangement considered in this work. This can potentially enhance the expressibility of the parameterized circuit while still avoiding the barren plateau.
A related important open question concerns the classical simulability of the proposed ansatze. Previous work [22] suggests that for qubits in 2D geometries, the finite local depth circuits adapted here are not classically efficiently simulatable. Opposite opinions exist arguing that any barren plateau-free variational ansatze must be classically simulatable [37]. We stress that our approach offers quantum advantages in the latter case as well: the training process can then be performed on a classical computer, outputting the optimal global gate sequences to be implemented on quantum hardware. The latter, then, for other purposes, prepares the target quantum state, which is nevertheless not fully accessible to classical computers.
Acknowledgment.—The authors thank Marco Cerezo, Martin Larocca, and Sabre Kais for helpful discussions. This work was supported in part by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, through the Quantum Internet to Accelerate Scientific Discovery Program, and in part by the LDRD program at Los Alamos. V.S. also acknowledges support from the Quantum Computing Summer School at Los Alamos National Laboratory, and support from the U.S. Department of Energy Quantum Science Center.

### v2

@@ Abstract
In conventional circuit-based quantum computing architectures, the standard gate set includes arbitrary single-qubit rotations and two-qubit entangling gates. This choice is not always aligned with the native operations available in certain hardware, where the natural entangling gates are not restricted to two qubits but can act on multiple, or even all, qubits simultaneously. However, leveraging the capabilities of global quantum operations for algorithm implementations is highly challenging, as directly compiling local gate sequences into global gates usually gives rise to a quantum circuit that is more complex than the original one. Here, we circumvent this difficulty using a variational approach. Specifically, we study parameterized circuit ansatze composed of a finite number of global gates and layers of single-qubit unitaries. We demonstrate the expressibility of these ansatze and apply them to the problem of ground state preparation for the Heisenberg model and the toric code Hamiltonian, highlighting their potential for offering practical advantages.
@@
In this work, we address this gap with a variational approach (Figure 1). Specifically, we implement variational quantum algorithms using circuit ansatzes that are directly constructed from alternating layers of global gates and single-qubit rotations. This design restricts the circuit to a constant number of global gates, ensuring that the implementations remain low-cost and are suitable for near-term quantum computers supporting such gates. Despite these constraints, we demonstrate that the ansatzes are highly expressive and capable of preparing a diverse range of quantum states, including those with long-range entanglement.
@@ I Global Gates
Global gates are unitary operations that simultaneously act on multiple or all qubits in a system. They are native to several platforms and have been demonstrated experimentally across a range of hardware [9, 10, 11, 5, 12, 13, 14, 16, 17, 18, 19, 20, 21]. They are typically achieved by driving the systems through their native underlying interactions and with global control pulses. For example, in Rydberg atom systems, global operations can be implemented using resonant laser pulses over a static atomic arrangement [14]. Superconducting systems can emulate global gates by coupling the Josephson junction qubits to a shared resonator mode [22].
@@
In addition, as single-qubit gates can be implemented efficiently in high fidelity with a small execution time compared to the entangling operations on the physical device, we treat single-qubit gates as a free resource. This also makes working with a variety of entangling gates equivalent so long as they can be obtained from the GCZ gates using a simple single-qubit basis change. For instance, we treat the GCX gate
@@ II.1 Barren Plateau
One key requirement for any successful ansatz is its ability to train well, starting from any arbitrary initial state and for a wide variety of Hamiltonians. The gradient-based optimizers rely on sufficiently large gradient values so that the parameters of the ansatz can be updated. However, it is well known that sufficiently complex ansatze (as random as a unitary $2$-design) admits the infamous “barren plateau” phenomenon [26, 27]. That is the variance of the gradient of the cost function, ${\rm Var}\left[\partial_{\vec{\theta}}C\right]$ decays exponentially with the system size, making the landscape of the cost function essentially flat and untrainable.
Numerous methods have been adopted to surpass the barren plateau issue, with various benefits and drawbacks. Here, we focus on a restricted class of circuits with depths no more than the logarithm of the number of qubits. Such circuits are known to be free of barren plateau [28, 29] when the input states and final measurements are further restricted to product state and local observables, respectively. However, such circuits can be efficiently simulated classically and hence do not directly offer quantum advantages. Nevertheless, the post-trained ansatze can used as an intermediate building block for subsequential applications. The crucial question asked is whether such circuit ansatze are sufficiently expressive. This question does not have a general answer and requires case-by-case analysis [27]. In the following, we will demonstrate the global-gate ansatze is sufficiently expressive to represent even long-range entangled states.
@@ II.2 Global Gate Ansatze
Our general strategy for constructing the global gate ansatze starts from building the circuit with a finite number ($k$) of alternating layers of single-qubit gates and two-qubit gates. See Figure 2 for an illustration. The two-qubit layers are designed to have finite local depth and are, therefore, guaranteed to be free of barren plateau. Crucially, each layer of two-qubit gates consists of only the same type of gates, i.e., ${\rm CZ}(\theta)$ as defined in (2), such that they can be implemented using a single global gate GCZ in equation (1).
We compare a variety of ansatze as constructed with the above general procedure. The detailed structures of the ansatze also depend on the underlying architecture (qubit connectivity). For simplicity, we first describe ansatze with qubits arranged in a $1$-D geometry. In this case, each layer of 2-qubit unitaries connects only neighboring sites in a ladder format (Figure 2). We also restrict our discussion to finite-depth ansatze, though they can be easily generalized to $\mathcal{O}(\log(n))$ depth. Our numerical simulations are limited to relatively small system sizes and, therefore, cannot reveal clear distinctions between finite and log-depth in both the circuit structures and the final performance in state preparations.
@@
  • Cartan Ansatz: We also compare the above ansatzes, which can be implemented with a finite number of global gates, to the most general circuits with local finite depth [29], where each two-qubit gate can be an arbitrary two-qubit unitary. The name is inspired from the Cartan decomposition of arbitrary 2-qubit unitaries. This ansatz cannot be efficiently realized using global gates but nevertheless serves as a benchmark for our global gate circuits.
The ansatze constructed using the above procedure has finite depth and is therefore barren plateau free. We illustrate this fact with numerical simulations in Figure 3 (a,b).
@@ IV Discussion
In this work, we developed a variational approach to harness the power of global quantum gates, native to many quantum computing platforms. Our proposed circuit ansatze consists of a finite number of global quantum gates interspersed with single-qubit rotations, enabling efficient implementation. We demonstrated the efficacy of this approach by using the global gate ansatze to prepare the ground states of the Heisenberg model and the Toric code model, achieving fast convergence and high accuracy.
Our approach is not limited to the global control-Z gate as considered, but can be naturally generalized to other sets of global operations. Another promising direction is to investigate whether one can use a finite (or logarithm) number of global gates to simulate circuits that, with respect to local quantum gates, have logarithmic local depth but total linear depth. Such circuits are also barren plateau free but can potentially offer direct quantum advantage.
Although the scenario we studied, i.e., finite-depth (shallow) circuits with local measurements, can be efficiently classically simulated and hence do not offer direct quantum advantages for statistics of local measurements alone, the post-training circuits can be ported as building blocks for subsequential tasks, for instance, preparing the physical target quantum states whose non-local correlations cannot be efficiently simulated. The circuit ansatze can be extended to other applications, such as unitary compiling or Hamiltonian discrimination using quantum machine learning [44]. It is crucial to emphasize that the high expressibility of the proposed global-gate ansatzes are irrespective of their classical simulatability with respect to local observables. Since shallow circuits, in general, do admit quantum advantages [45, 46, 47, 48, 49], it is interesting to investigate whether the proposed ansatze can be adapted to these scenarios as well.
Acknowledgment.—The authors thank Marco Cerezo, Martin Larocca, Manas Sajjan, and Sabre Kais for helpful discussions. This work was supported in part by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, through the Quantum Internet to Accelerate Scientific Discovery Program, and in part by the LDRD program at Los Alamos. V.S. also acknowledges support from the Quantum Computing Summer School at Los Alamos National Laboratory, and support from the U.S. Department of Energy Quantum Science Center.